# Non-Gaussian state teleportation with a nonlinear feedforward

### v4

@@ 3.1 Deterministic scheme
The nonlinear squeezing that can be achieved with a pure two-mode cluster state for the individual schemes is shown in Fig. 2 as a function of the maximal available Gaussian squeezing, see appendix B for details. The cluster state is optimized for each scheme with upper bound on Gaussian squeezing and the parameters of the optical circuit are searched to maximize the nonlinear squeezing at the output. Optimized preparation of the cubic state with finite nonlinear squeezing is subject to the same constraints. We also optimize the coefficients in superpositions (20) and (21) for best performance, thus obtaining states different to the eigenstates of restriction of the $O$ operator on finite Fock subspace.
The plot can be interpreted pairwise, according to the properties of used teleported state and its $N$-photon non-Gaussianity. This point of view connects together regular teleportation of the state (20) consisting of superposition of vacuum and single photon, and the nonlinear teleportation of the same state (albeit with different optimal coefficient $u$). Similarly we can compare the regular and nonlinear teleportation for ancillary state (21) and at last the schemes utilizing finitely squeezed cubic state (18). Such approach connects schemes possessing the same resources of non-Gaussianity and Gaussian squeezing with differences only in the classical part that processes measurement results and governs feedforwarded Gaussian operations.
@@
Teleportation of the weak nonlinear effects provided by the two component ancillary state (20) shows the strength of linear Gaussian processing as the nonlinear feedforward introduces no significant improvement in deterministic regime. When increasing the resource to two-photon non-Gaussianity (21) there is a slight difference for a realistic Gaussian squeezing of 6 and 7 dB available at the cluster state preparation. A major improvement by nonlinear feedforward appears for small Gaussian squeezing when considering finitely nonlinearly squeezed cubic state (18). The difference is even more pronounced for higher upperbound for Gaussian squeezing during optimization. Contrary, the regular teleportation does not gain much improvement with higher n-photon non-Gaussianity of the teleported state in the regime with 6 dB cluster state. The ideal nonlinear cubic phase measurement surpasses performance of all considered schemes. Altogether, when increasing the n-photon non-Gaussianity of the supporting states, the nonlinear feedforward starts manifesting advantage over teleportation, which lies in higher nonlinear squeezing at the output.
@@
Figure 2: Nonlinear squeezing at the output mode, when the available Gaussian squeezing is limited to certain value [dB] during optimization. The two-mode cluster state is pure and optimized for each scheme. Output nonlinear squeezing is shown for regular and nonlinear teleportation of the superposition of vacuum state and single photon (20) and the state with 2-photon non-Gaussianity (21), both with optimized coefficients. Finally the results obtained with the finitely squeezed cubic state (18) and nonlinear cubic phase measurement.
@@
Figure 3: Nonlinear squeezing at the output mode, when the available Gaussian squeezing is limited to certain value $[dB]$ during optimization. The two-mode cluster state carries $n=0.1$ additional thermal noise.
@@
Fig. 3 shows nonlinear squeezing transferred via the schemes when the cluster state is initially prepared not from a pure vacuum state but from thermal states that carry $n=0.1$ additional vacuum fluctuations. The following optimization of the cluster state and the schemes is as for the previous plot within bounded Gaussian squeezing. The performance for all schemes is negatively affected, however their relative comparison remains.
@@ 3.2 Probabilistic regime
Let us start with more detailed discussion of some cases presented in Fig. 2. To fully understand the role of the nonlinear feedforward, we will look at teleportation via 6 dB, 7 dB and 8 dB cluster state in a probabilistic regime, at all the measurement outcomes individually. We choose the two component superposition (20) for this analysis as it is close to experimental reality and the parameters of optical schemes and input states are taken from the optimized deterministic regime, thus they are the same as for which the results shown in Fig. 2 were obtained.
As the output mixture cover more states, the overall probability of success or obtaining such state increases, whereas the nonlinear squeezing diminishes, as can be seen in Fig. 4. This happens as a result of accepting less nonlinearly squeezed states to the mixture. As we increase the squeezing of cluster state, the advantage brought by the nonlinear feedforward vanishes. To see the behaviour as we decrease the cluster state squeezing we added a case of 4dB and 5dB cluster states, however it seems that from discussed cases, teleportation through the 6dB cluster state manifests the largest difference provided by the nonlinear feedforward.
@@
We conclude our analysis by comparing performance of unity gain teleportation with and without the nonlinear feedforward. As the teleported state is considered the experimentally realistic two component superposition (20). Its specific form given by $u=0.78977467$ is chosen such that the state is eigenstate of the operator $O$ (5) restricted in Fock space to the subspace of dimension two. This approach ensures maximal nonlinear squeezing that can be obtained with superposition of vacuum and single photon state. The nonlinear squeezing is evaluated at this states’ native cubicity. Thus a state with maximal nonlinear squeezing on a Fock subspace of dimension two is teleported and aftewards evaluation is done at its native cubicity before the teleportation. This setting leaves us with a single degree of freedom, $\chi$ in the nonlinear feedforward, which is optimized and yields $\chi=-0.219$ in the case of 6dB cluster state and $\chi=-0.130$ for 8dB cluster state.
Figure 5: Scheme for conditioned regular/nonlinear teleportation including the preparation of the ancillary state.
Nonlinear squeezing shown in Fig. 6 is visibly enhanced by nonlinear feedforward and enables its detection even for probabilities for which the regular teleportation fails to transfer any. We also introduced $25\%$ losses to the ancillary state, that was present in its recent experimental preparation. The results are shown in dashed lines and again show advantage of the nonlinear teleportation with nonlinear feedforward. The optimal strength of the nonlinear feedforward is higher than for the ideal case, yielding $\chi=-0.250$ for 6dB cluster state and $\chi=-0.151$ for 8dB cluster state. A similar effect was recognized in decoherence of the nonlinear squeezing [Kala22], where the native cubicity of the state increases with higher losses. In the presence of losses, the performances of teleportations through cluster states with 6 and 8 dB of Gaussian squeezing does not converge even for postselection on the best cases with vanishing probability.
Figure 6: Nonlinear squeezing (16) produced by unity gain regular/nonlinear teleportation of an eigenstate of the operator $O$ (5) conditioned on measurement results of the homodyne measurement with best results aggregated up to some probability and evaluated at native cubicity of the state before teleportation. The nonlinear teleportation and regular teleportation was simulated with two mode squeezed vacuum of 6 dB and 8 dB with (dashed lines) and without losses on the teleported state. Here it is assumed $25\%$ losses [Konno2021]. In all cases the nonlinear squeezing firstly vanishes (the plot line crosses the dashed line of boundary 1) for regular teleportation.
@@ 4 Conclusion
We showed an improvement in the teleported nonlinear squeezing by involving currently experimentally tested nonlinear feedforward [sakaguchi]. In a deterministic regime, the positive effect can be seen on the theoretical level of teleportation of the ideal cubic state. However in a probabilistic regime, when best results are postselected up to some probability, the improvement is observable even with current experimental resources in comparison to unity gain teleportation. Processing the measurement data during teleportation via classical nonlinear feedforward decreases the added amount of noise which can be seen by improved nonlinear squeezing at the output. The cubic nonlinear squeezing is a resource for measurement-based quantum computation [Konno2021, sakaguchi] and improving its amount that is transferred via teleportation increases the potential to chain and iterate computational steps as required by cluster computing.
The cubic nonlinearity itself is not currently experimentally available in travelling light, however a superposition of Fock states with nonlinear squeezing has already been prepared [Konno2021] as an alternative resource substituting the nonlinearity. Secondly, we show that in probabilistic regime the nonlinear teleportation equipped with nonlinear feedforward and currently experimentaly accessible state outperforms the unity gain linear teleportation through 6dB cluster state as is shown in Fig. 6. The advantage is thus in principle observable even with currently available nonlinearly squeezed state and appears by adjusting the classical part of the scheme consisting of nonlinear feedforward combined with post-selection. The enhancement remains observable even in the presence of losses that accompanied the experimental preparation [Konno2021]. Our work is one of the first steps required to extend the presently Gaussian theory of cluster states into its final destination - non-Gaussian domain.
@@ Acknowledgement
VK and PM acknowledge grant 22-08772S and RF grant 21-13491X of Czech Science Foundation. VK acknowledges the Quantera project CLUSSTAR (8C24003) of MEYS, Czech Republic. Project CLUSSTAR has received funding from the European Union’s Horizon 2020 Research and Innovation Programme under Grant Agreements No. 731473 and No. 101017733 (QuantERA). PM acknowledges a grant from the Programme Johannes Amos Comenius under the Ministry of Education, Youth and Sports of the Czech Republic reg. no. CZ.02.01.01/00/22_008/0004649. VK, PM and RF also acknolwedge Horizon Europe Research and Innovation Actions under Grant Agreement no. 101080173 (CLUSTEC). P.M. and R.F. acknowledge project CZ.02.01.010022 0080004649 (QUEENTEC) of EU and the Czech Ministry of Education, Youth and Sport. M.W. acknowledges funding from the Plan France 2030 through the project ANR-23-PETQ-0013. VK acknowledges project IGA-PrF-2024-008.
@@ Appendix A Model in Heisenberg picture
The schemes of measurement-induced nonlinearity and teleportation can be described by one model as they differ only by feedforwarded rotation parameterized by $\chi$ which can be switched of in the case of teleportation by setting $\chi=0$. The measurement results of $x$-homodyne and p-homodyne in b) and c) of Fig. 1 yields
@@
where quadratures with subscript $Q$ are of the ancillary mode and with $A$ of the first mode of the cluster state; $\theta$ depends on the measurement outcome of $x$-homodyne as
@@
The Gaussianity of the cluster state can be used for expressing the whole expression in terms appearing in covariance matrix of the modes A and B. For the more complex moments involving more than two quadratures, we use the generalized Stein’s lemma for multivariate Gaussian distributions of a random variable $\vec{X}$ with mean values $\vec{\mu}$ [Liu1994]

### v5

@@ 3.1 Deterministic scheme
Fig. 2 shows the nonlinear squeezing that can be transferred via an optimized pure two-mode cluster state ($n=0$) for the individual schemes Fig.1b, c, and d. It is plotted as a function of the maximal Gaussian squeezing available for the cluster state preparation and preparation of the state (6), see Appendix B for details. Within this limit, the squeezing $r_{1}$ and $r_{2}$ forming the cluster state is optimized together with its other parameters $\phi,\phi_{1},\phi_{2},\alpha_{1}$, and $\alpha_{2}$. The cluster state is optimized for each scheme to transfer the maximal amount of the nonlinear squeezing. The teleported or auxiliary states are optimized as well; for example, we optimize the coefficients in superpositions (8) and (9) for best performance, thus obtaining states different from eigenstates of the restriction of the $O$ operator on finite Fock subspace.
Fig. 2a can be interpreted pairwise, according to the properties of the used teleported/auxiliary state and its 11- or $2$-photon non-Gaussianity. The state (8), formed by a superposition of vacuum and the single photon state, has genuine $1$-photon non-Gaussianity. We can compare the performance of the canonical teleportation protocol with linear feedforward Fig.1c and the nonlinear teleportation Fig.1d when this superposition is used as the teleported state. Thus, the two schemes are initially equipped with a resource of non-Gaussianity of the same level. The teleported state is optimized individually, yielding different $u$ for different schemes (and thus also different initial nonlinear squeezing). Similarly, we can compare the canonical and nonlinear teleportation for the teleported state (9) and, at last, the schemes utilizing a finitely squeezed cubic state (6). In such a pairwise comparison, the initial level of non-Gaussianity is equal, and the only difference is whether the feedforward is linear or nonlinear. Generally, higher squeezing of the cluster state improves the transferred nonlinear squeezing. Fig. 2b shows the initial nonlinear squeezing of teleported/auxiliary states.
@@
The significant advantage of nonlinear feedforward manifested with the cubic state (6) nearly vanishes for the two-component superposition (8). This shows the relative strength of linear Gaussian processing of weak non-Gaussian states in the deterministic regime. When increasing the resource to two-photon non-Gaussianity (9) there is a slight difference for a realistic maximal Gaussian squeezing of 6 and 7 dB of the cluster state. A major improvement by nonlinear feedforward appears for small Gaussian squeezing when considering finitely nonlinearly squeezed cubic state (6). The difference is even more pronounced for higher maximal cluster state squeezing. The ideal nonlinear cubic phase measurement, Fig. 1b, surpasses the performance of all considered schemes. For comparison, a 10dB cluster state transfers -1.9 dB of nonlinear squeezing via the regular teleportation (Fig. 1c) and -2.9 dB via the nonlinear teleportation (Fig. 1d) of the finitely squeezed cubic state and -3.7 dB with the ideal projection (Fig. 1b). Interestingly, the slope with which the nonlinear squeezing improves with increasing cluster state squeezing is comparable for all three schemes, when the cubic state is involved, or projected on. Altogether, when increasing to $2$-photon non-Gaussianity of the supporting state (9), the nonlinear teleportation with nonlinear feedforward starts manifesting an advantage over teleportation, which lies in higher nonlinear squeezing at the output.
@@
Figure 2: a) Nonlinear squeezing at the output mode, expressed in dB (3). The Gaussian squeezing, available for the preparation of the cluster state and the state (6), is limited to a certain value [dB] during optimization. The limit is shown on the horizontal axis. The two-mode cluster state is pure and optimized for each scheme. Output nonlinear squeezing is shown for deterministic regular and nonlinear teleportation (Fig. 1c and d) of the superposition of vacuum state and single photon (8) and the state with 2-photon non-Gaussianity (9), both with optimized coefficients. Finally, results are presented for the finitely squeezed cubic state (6) serving as teleported/ auxiliary state and nonlinear cubic phase measurement shown in Fig. 1b. b) The subplot shows the initial nonlinear squeezing before teleportation/measurement.
@@
Figure 3: a) Nonlinear squeezing at the output mode, expressed in dB (3). The Gaussian squeezing, available for the preparation of the cluster state and the state (6), is limited to a certain value [dB] during optimization. The limit is shown on the horizontal axis. The two-mode cluster state carries $n=0.1$ additional thermal noise. To better see the comparison with the case of a pure cluster state (Fig. 2a), the light gray shadow shows results for $n=0$. b) The subplot shows the initial nonlinear squeezing before teleportation/measurement.
@@
Fig. 3a shows nonlinear squeezing transferred via the schemes in Fig. 1b,c, and d when the cluster state is initially prepared not from a pure vacuum state but from thermal states that carry $n=0.1$ additional fluctuations over the vacuum. The following optimization of the cluster state and the schemes is as for the previous plot within bounded Gaussian squeezing. The performance of all schemes is negatively affected; the nonlinear squeezing transferred with the pure cluster state is depicted with the light shadow. The relative comparison of the schemes remains unchanged. Fig. 3b shows the initial nonlinear squeezing.
@@ 3.2 Probabilistic regime
In the previous section, we showed the potential of the nonlinear feedforward. In the deterministic regime, the advantage is more pronounced for the experimentally challenging cubic state. Here, we will turn our attention to the probabilistic regime, seeking a case that enables showing the advantage of the nonlinear feedforward even with currently accessible states. Let us start with a more detailed discussion of some cases presented in Fig. 2 a. To fully understand the role of the nonlinear feedforward, we will look at teleportation via 6 dB, 7 dB, and 8 dB cluster states in a probabilistic regime at all the measurement outcomes individually. We choose the two-component superposition (8) for this analysis as it is close to experimental reality, and the parameters of optical schemes and input states are taken from the optimized deterministic regime. Thus they are the same as for which the results shown in Fig. 2 were obtained.
As more measurement results are accepted and therefore, the output mixture includes more states, the overall probability of success or obtaining such a state increases, whereas the nonlinear squeezing diminishes, as can be seen in Fig. 4. This happens as a result of accepting less nonlinearly squeezed states to the mixture. As we increase the squeezing of the cluster state, the advantage brought by the nonlinear feedforward vanishes. To see the behavior as we decrease the cluster state squeezing, we added a case of 4dB and 5dB of squeezing in the cluster state; however, from discussed cases, teleportation through the 6dB cluster state manifests the largest difference provided by the nonlinear feedforward.
Measurement results around the origin of the phase space alter the teleported state only slightly, which diminishes the difference between the linear and nonlinear feedforward in schemes Fig. 1c and d. Increasing the probability means including larger measurement results. At a certain point, a discrepancy appears between the nonlinear feedforward designed for the ideal cubic state and the auxiliary state, which is the two-component superposition. This leads to similar performance of the canonical teleportation and nonlinear measurement at higher probabilities. The difference is best seen for the success probability $0.2$ in Fig. 4.
@@
We conclude our analysis by comparing the performance of unity gain teleportation with and without the nonlinear feedforward , shown in Fig 5. As the teleported state is considered the experimentally realistic two-component superposition (8). Its specific form given by $u=0.79$ is chosen such that the state is an eigenstate of the operator $O$ in Eq. (1) restricted in Fock space to the subspace of dimension two. This approach ensures maximal nonlinear squeezing that can be obtained with the superposition of vacuum and single-photon state. The nonlinear squeezing is evaluated at this state’s native cubicity $z_{n}$ (10). Thus, a state with maximal nonlinear squeezing on a Fock subspace of dimension two (8) is teleported, and afterward, evaluation is done at its native cubicity before the teleportation. This setting leaves us with a single degree of freedom, $\chi$ in the nonlinear feedforward, which is optimized and yields $\chi=-0.219$ in the case of 6dB cluster state and $\chi=-0.130$ for 8dB cluster state.
Figure 5: Scheme for conditioned regular/nonlinear teleportation with the two-mode squeezed vacuum including the preparation of the teleported state. The outcoming state is accepted or discarded based on the measurement results $m_{x}$ and $m_{p}$.
Nonlinear squeezing shown in Fig. 6 is visibly enhanced by nonlinear feedforward and enables its detection even for probabilities for which the regular teleportation fails to transfer any. We also introduced $25\%$ losses to the auxiliary state, which was present in its recent experimental preparation. The results are shown in dashed lines and again show the advantage of the nonlinear teleportation with nonlinear feedforward. The optimal strength of the nonlinear feedforward is higher than for the ideal case, yielding $\chi=-0.250$ for 6dB cluster state and $\chi=-0.151$ for 8dB cluster state. A similar effect was recognized in the decoherence of the nonlinear squeezing [21], where the native cubicity of the state increases with higher losses. In the presence of losses, the performances of teleportations through cluster states with 6 and 8 dB of Gaussian squeezing do not converge even for postselection on the best cases with vanishing probability.
Figure 6: Nonlinear squeezing (23) produced by unity gain regular/nonlinear teleportation of an eigenstate of the operator in Eq. $O$ (1) conditioned on measurement results of the homodyne measurement with best results aggregated up to some probability and evaluated at native cubicity of the state before teleportation. The nonlinear teleportation and regular teleportation were simulated with two mode squeezed vacuum of 6 dB and 8 dB with (dashed lines) and without losses on the teleported state. Here it is assumed $25\%$ losses [24]. In all cases the nonlinear squeezing vanishes first (the plot line crosses the dashed line of boundary 1) for regular teleportation.
@@ 4 Conclusion
Instead of a teleportation of the typical Fock states [37] or cat states [39], here, we consider cubic nonlinear squeezing, a notion of purely continuous-variable non-Gaussianity, that has a straightforward interpretation as a resource for nonlinear quantum operations [29, 7]. We showed an improvement in the teleported nonlinear squeezing by involving currently experimentally tested nonlinear feedforward [43]. In a deterministic regime, the positive effect can be seen on the theoretical level of teleportation of the ideal cubic state. However in a probabilistic regime, when best results are postselected up to some probability, the improvement is observable even with current experimental resources in comparison to unity gain teleportation. Processing the measurement data during teleportation via classical nonlinear feedforward decreases the added amount of noise which can be seen by improved nonlinear squeezing at the output. The cubic nonlinear squeezing is a resource for measurement-based quantum computation [24, 43] and improving how much of it is transferred via teleportation increases the potential to chain and iterate computational steps as required by cluster computing. The cubic nonlinearity itself is not currently experimentally available in traveling light. However, a superposition of Fock states with nonlinear squeezing has already been prepared [24] as an alternative resource substituting the nonlinearity and a design based on the Kerr nonlinearity is also proposed for near future tests [54].
Secondly, we show that in a probabilistic regime, the nonlinear teleportation equipped with nonlinear feedforward and currently experimentally accessible state outperforms the unity gain linear teleportation through 6dB cluster state as is shown in Fig. 6. The advantage is thus, in principle, observable even with the currently available nonlinearly squeezed state and appears by adjusting the classical part of the scheme consisting of nonlinear feedforward combined with post-selection. The enhancement remains observable even in the presence of losses that accompanied the experimental preparation [24]. Our work is one of the first steps required to extend the presently Gaussian theory of cluster states into its final destination - the non-Gaussian domain.
@@ Acknowledgement
VK and PM acknowledge grant 22-08772S and RF grant 21-13491X of Czech Science Foundation. VK acknowledges the Quantera project CLUSSTAR (8C24003) of MEYS, Czech Republic. Project CLUSSTAR has received funding from the European Union’s Horizon 2020 Research and Innovation Programme under Grant Agreements No. 731473 and No. 101017733 (QuantERA). PM acknowledges a grant from the Programme Johannes Amos Comenius under the Ministry of Education, Youth and Sports of the Czech Republic reg. no. CZ.02.01.01/00/22_008/0004649. VK, PM and RF also acknolwedge Horizon Europe Research and Innovation Actions under Grant Agreement no. 101080173 (CLUSTEC). P.M. and R.F. acknowledge project CZ.02.01.010022 0080004649 (QUEENTEC) of EU and the Czech Ministry of Education, Youth and Sport. M.W. acknowledges funding from the Plan France 2030 through the project ANR-23-PETQ-0013. VK acknowledges project IGA-PrF-2025-008.
@@ Appendix A Model in Heisenberg picture
The schemes of measurement-induced nonlinearity and teleportation can be described by one model as they differ only by feedforwarded rotation parameterized by $\chi$ which can be switched of in the case of teleportation by setting $\chi=0$. The measurement results of $x$-homodyne and p-homodyne in c) and d) of Fig. 1 yields
@@
where quadratures with subscript $Q$ are of the auxiliary mode and with $A$ of the first mode of the cluster state; $\theta$ depends on the measurement outcome of $x$-homodyne as
@@
The Gaussianity of the cluster state can be used for expressing the whole expression in terms appearing in covariance matrix of the modes A and B. For the more complex moments involving more than two quadratures, we use the generalized Stein’s lemma for multivariate Gaussian distributions of a random variable $\vec{X}$ with mean values $\vec{\mu}$ [28]